OpenStudy (anonymous):

Expand the series and Evaluate: Enter your answer as the example: 1+2+3+4+5=15 sum_(k=1)^5(k^3)

OpenStudy (anonymous):

replace k by 1, then 2, then 3, then 4, then 5 put a plus sign between all the terms

OpenStudy (anonymous):

in other words it is \[\sum_{k=1}^5k^3=1^3+2^3+3^3+4^3+5^3\] then some arithmetic

OpenStudy (anonymous):

1+8+27+64+125=225

OpenStudy (anonymous):

that'll do it. you can also use \[\sum_{k=1}^nk^3=\left(\frac{n(n+1)}{2}\right)^2\]